OpenStudy (anonymous):

Three coins are tossed in succession. What is the probability of heads-tails-heads? 1/2 1/6 1/8

OpenStudy (anonymous):

1/6 there are 2 sides to each coin :)

OpenStudy (anonymous):

okay fine, 3 coins two sides each. = 6 sides. you have 1 chance of getting one of those sides. =1/6

Parth (parthkohli):

The probability of a head is 1/2, and the probability of a tail is 1/2. You multiply the probabilities of each event to get the new probability :)

Parth (parthkohli):

\( \color{Black}{\Rightarrow\Large {1 \over 2} \cdot {1 \over 2} \cdot {1 \over 2} = {1 \over 2^3} }\)

OpenStudy (anonymous):

oh. .-. oops

Parth (parthkohli):

Nah it's all fine

OpenStudy (anonymous):

so its 1/8

OpenStudy (anonymous):

yeah. Parthicus was right